OpenStudy (anonymous):

I NEED HELP PLEASE! Solve the following simultaneous equations using the Elimination method: 2x-3y=1 11y-7x=5 PLEASE HELP!

hero (hero):

Isolate x in the first equation, then substitute the corresponding expression into the second equation

OpenStudy (anonymous):

Can you write the whole thing for me? please? coz tomorrow i have to submit this worksheet.

jhonyy9 (jhonyy9):

multiplie first equation by 7 and multiplie second equation by 2 and assum 14x -21y =7 -14x +22y=10 ------------------ 0 y=17 2x-3(17)=1 2x-51=1 2x=52 x=52/2 x=26 x=26 y=17